OpenStudy (anonymous):

fractional notation 62.5%

OpenStudy (anonymous):

this is the last one guys

OpenStudy (cathyangs):

5/8

OpenStudy (anonymous):

625/100 = 125/20 = 25/4

OpenStudy (anonymous):

so is it 5/8 or 25/4

OpenStudy (anonymous):

no its 5/8, I made a mistake

OpenStudy (anonymous):

sorry

OpenStudy (cathyangs):

um...well. 62.5% of something is a little over a half. it's writtten in decimal form as .625...

OpenStudy (anonymous):

oh ok i thank you guys so much u all been excellent help thank u

OpenStudy (cathyangs):

welcome!